OpenStudy (anonymous):

what is the 20th term of the sequence that begins -3,6,-12,24 A.-524,288 B.524,288 C.1,572,864 D.3,145,728

OpenStudy (raden):

looks this is a geometri sequence, right ?

OpenStudy (raden):

with a1 = -3 and ration (r) = -2

OpenStudy (raden):

use the formula ; an = a1 * r^(n-1)

OpenStudy (raden):

so, a20 = a1 * r^(20 - 1) = a1 * r^19

OpenStudy (anonymous):

so is it 1,572,864?, thats what i got when i used that formula

OpenStudy (raden):

yes, i got like yours good job :)

OpenStudy (anonymous):

thank you for the help!

OpenStudy (raden):

wlcm